Find the Mean value of the function f left parenthesis x right parenthesis space equals space x cubed space minus space 3 x squared over the interval [0, 5].

1
Expert's answer
2021-11-08T06:11:03-0500

Let us find the Mean value of the function "f(x)=x^3 - 3 x^2" over the interval "[0, 5]."

It follows that the Mean value of the function is

"\overline{f}=\frac{1}{5-0}\int\limits_0^5f(x)dx
=\frac{1}{5}\int\limits_0^5f(x^3-3x^2)dx
=\frac{1}{5}(\frac{x^4}4-x^3)|_0^5\\
=\frac{1}5(\frac{625}4-125)=\frac{1}5\cdot\frac{125}4=\frac{25}{4}=6.25"